OpenStudy (anonymous):

find the volume of the cylinder to the nearest cubic foot. use a calculator WILL GIVE MEDAL FOR CORRECT ANSWER! Read comments section for link

OpenStudy (anonymous):

c'mon people!

OpenStudy (mathstudent55):

The volume of a cylinder is: \(V = \pi r^2h\) where r= radius of the base, and h = height of the cylinder. For r = 2 ft, and h = 2 ft, you have: \(V = \pi(2 ~ft)^2(2 ~ft) \) \(V = \pi (4 ft^2)(2 ft) \) \(V = 8\pi ft^3 \) \(V = 25.13 ft^3 \) Round off to the nearest cubic foot: \(V = 25 ft^3 \)

OpenStudy (anonymous):

Thanks

OpenStudy (mathstudent55):

You're welcome.